# Generalizable and Relightable Gaussian Splatting for Human Novel View Synthesis

### v1

@@ 3.2 Physically Grounded Neural Rendering
Based on the lighting-invariant geometry, we further propose a PGNR module to combine the generalization capabilities of neural networks with the physical accuracy of physically based rendering, enabling efficient inference of material properties and complex illumination of unseen data.
@@
To support relightable rendering, PGNR parameterizes each 3D Gaussian point $\mathcal{P}$ as a set of attributes: (1) basic attributes: position $\mathbf{p}\in\mathbb{R}^{3}$, rotation $\mathbf{r}\in\mathbb{R}^{4}$, scale $\mathbf{s}\in\mathbb{R}^{3}$, and opacity $\alpha\in\mathbb{R}$, (2) geometry-related attributes: surface normal $\mathbf{n}\in\mathbb{R}^{3}$ and light visibility $\mathbf{v}\in\mathbb{R}^{16}$ encoded via SH coefficients, (3) material attributes: albedo $\mathbf{a}\in\mathbb{R}^{3}$ and roughness $\gamma\in\mathbb{R}$, and (4) illumination attributes: indirect lighting $\mathbf{l}_{\text{ind}}\in\mathbb{R}^{48}$ represented by SH coefficients. We formulate these 3D attributes on corresponding 2D maps via pixel-aligned depth maps (Sec. 3.1), allowing direct supervision in image space. For simplicity, we consider a single-view depth map $\mathbf{D}$ and compute 3D position map $\mathbf{X}$ via unprojection (Eq. 3), while normals $\mathbf{N}_{f}$ are predicted by the refinement module. For other Gaussian attribute maps, we introduce a geometry-aware encoder $\xi_{\text{geo}}$ that extracts multi-scale geometric features $\{\mathbf{F}_{s}^{g}\}_{s=1}^{K}$ from $\mathbf{D}$ and $\mathbf{N}_{f}$, encoding both 2D image features and 3D spatial geometry. These features are then fused with radiance features to incorporate appearance and geometric context for jointly modeling geometry-aware Gaussian attributes. Besides, since the SH-encoded visibility map $\mathbf{Y}_{\text{vis}}$ is independent of image appearance, we predict it via another lightweight decoder $\Phi_{\text{vis}}$ conditioned solely on the geometric features $\{\mathbf{F}_{s}^{g}\}_{s=1}^{K}$. As for the remaining Gaussian maps, we extract the full-resolution Gaussian features $\Theta$ through a decoder $\Phi_{\text{parm}}$:
@@
where $\oplus$ denotes the concatenation operation applied across all feature levels. Subsequently, the vanilla Gaussian parameter maps ($\mathbf{R}$, $\mathbf{S}$, and $\mathbf{O}$) are predicted with a Gaussian head $h_{g}$:
@@
To accelerate the convergence of albedo estimation under diverse lighting conditions, we introduce a residual component $\Delta\mathbf{A}$, which serves as a delighting term to better disentangle shadows and illumination effects from intrinsic surface properties. We achieve this by leveraging a material prediction head $h_{m}$ to jointly estimate the residual albedo $\Delta\mathbf{A}$ and the surface roughness map $\mathbf{\Gamma}$:
@@
The final albedo is computed as: $\mathbf{A}=\text{Sigmoid}\left(\mathbf{I}+\Delta\mathbf{A}\right)$, where $\mathbf{I}$ denotes the input image.
@@
[…]
where $\bm{\omega}$ and $\bm{\omega}^{\prime}$ denote spherical directions in the original and convolved environment map respectively, $\Omega$ represents the unit sphere, and $l$ is the shininess exponent controlling the angular falloff in the Phong reflectance model. However, such integration may result in underexposed lighting compared to ground-truth shading, as illustrated in Sec. 4.2. To address this, we introduce a direct illumination scaling factor $s_{\text{d}}\in\mathbb{R}$ to globally compensate for the brightness discrepancy. In addition, since the SH-encoded indirect light map $\mathbf{Y}_{\text{ind}}$ is influenced not only by the appearance and geometry but also by the input illumination, we design a light encoder-decoder that predicts the direct illumination scaling map $\mathbf{S}_{\text{d}}$ and the indirect light map $\mathbf{Y}_{\text{ind}}$:
@@
To enable physically plausible light interaction on the human surface, we compute the PBR color $\mathbf{C}_{\text{PBR}}$ for each Gaussian point. Following the rendering equation [19], the outgoing radiance $\mathbf{C}_{\text{PBR}}$ in direction $\bm{\omega}_{o}$ can is given by:
[…]
where $f$ is the simplified Disney BRDF function [5] modeling the surface reflectance properties, and $\Omega$ here represents the hemisphere oriented around the surface normal $\bm{n}$. $\mathbf{L}\left(\bm{\omega}_{i}\right)$ denotes the incident radiance from direction $\bm{\omega}_{i}$:
$\mathbf{L}\left(\bm{\omega}_{i}\right)=V\left(\bm{\omega}_{i}\right)\left(s_{\text{d}}\cdot\mathbf{L}_{\text{d}}^{\prime}\left(\bm{\omega}_{i}\right)\right)+\left(1-V\left(\bm{\omega}_{i}\right)\right)\mathbf{L}_{\text{ind}}\left(\bm{\omega}_{i}\right)$ (14)
where $V\left(\bm{\omega}_{i}\right)$ denotes the light visibility from direction $\bm{\omega}_{i}$, parameterized using SH coefficients $\mathbf{v}$. Similarly, $\mathbf{L}_{\text{ind}}\left(\bm{\omega}_{i}\right)$ represents the indirect illumination from direction $\bm{\omega}_{i}$, also encoded via SH coefficients $\mathbf{l}_{\text{ind}}$. Although R3DGS [12] employs a similar lighting model, it estimates indirect illumination in an implicit manner through unsupervised inverse rendering, which often results in inaccurate predictions and limited generalization to novel lighting conditions while imposing a substantial computational burden due to the explicit ray tracing strategy. In contrast, our GRGS utilizes the strong generalization capability of a U-Net-based architecture in conjunction with a 2D-to-3D projection training strategy, enabling more accurate estimation of both visibility and indirect illumination under arbitrary lighting. After computing the physically based rendering color $\mathbf{C}_{\text{PBR}}$ for each Gaussian point, we perform rasterization to synthesize the final image.
@@
Leveraging the excellent differentiability of 3DGS [21], GRGS directly projects 2D image supervision into the 3D space to avoid heavy computation of explicit ray tracing. Specifically, ambient occlusion, direct light, and indirect light maps are used as photometric supervision signals and lifted into the Gaussian representation to enable efficient gradient-based optimization. To improve alignment between 2D projections and 3D geometry, we further enhance ambient occlusion supervision with direct light maps, which implicitly constrain visibility through shadow-aware regions. For better learning of indirect illumination, we employ a gradient-truncated hard shadow fusion scheme within the rendering equation, enabling the effective disentanglement of direct and indirect lighting components.
@@ 4 Experiments
We utilize two human scan datasets, Twindom [46] and THuman2.0 [57], to validate the effectiveness of our method. To ensure high-quality rendering, we carefully selected 800 scans from Twindom and 337 scans from THuman2.0, filtering out meshes with noticeable artifacts. Additionally, 384 HDR environment maps were collected from Polyhaven, HDRMAPS, and iHDRI to simulate diverse illumination conditions. More dataset construction details are listed in Supp.Mat..
Table 1: Quantitative comparison on the synthetic dataset. In accordance with [63], the SSIM and LPIPS metrics are calculated within the bounding box delineating the human region, while the MAE of normal and PSNR metrics is computed within the foreground mask.
@@
Figure 4: Qualitative comparison of our method and 2D image-based methods. Zoom in for the best view.
@@
We compare our method on our synthetic dataset and real dataset [63] against two 3DGS-based approaches, R3DGS [12] and ARGS [26], as well as two 2D image-based methods, IC-Light [60] and SwitchLight [22]. Note that since 2D methods cannot synthesize novel views, we directly use ground-truth novel view images as their input for a fair comparison. For quantitative evaluation, we employ PSNR, SSIM, and LPIPS to evaluate the diffuse albedo, ambient occlusion, and relighting quality. Further, we utilize MAE to evaluate the normal map.
@@ 4.1 Comparison Results
Figure 5: Geometry consistency comparison.
We first make a comparison with 3D-based methods (i.e., R3DGS [12] and ARGS [26]) from the perspective of material estimation, geometry reconstruction, and relighting quality, as shown in Fig. 3 and Table 1. R3DGS [12], as a general approach, struggles to recover smooth and detailed surface normals from sparse-view inputs due to the absence of human-specific geometric priors, resulting in suboptimal relighting performance. ARGS [26] mitigates this to some extent by introducing a body template prior, which, in contrast, yields overly smooth geometry. Thanks to the proposed LGR and PGNR modules, the proposed GRGS produces high-quality geometry and realistic relighting results under novel viewpoints.
Further, the comparisons with 2D-based methods (i.e., IC-Light [60] and SwitchLight [22]) are illustrated in Fig. 4. The baseline methods produce noticeable shading and lighting variations, which, however, lack physical plausibility, thereby undermining overall realism. In contrast, our results closely approximate the ground-truth images generated by a physically-based renderer and achieve superior perceptual quality. This is further corroborated by a user study (see Supp.Mat.), which confirms that our approach yields more visually plausible and realistic relighting results than baseline methods.
@@ 4.2 Ablation Study
Lighting-invariant Prior.
A major limitation of the original GPS-Gaussian [63] is that it struggles to maintain robustness against uneven lighting on the human surface without lighting priors, leading to a significant degradation in geometric consistency, as shown in Fig. 5. In contrast, our LGR module learns lighting-invariant radiance features prior under diverse lighting conditions and thus enhancing the reconstruction robustness to uneven illumination, enabling both geometric consistency and NVS quality improvements (Fig. 6) .
Normal Refinement.
Although the lighting-invariant prior improves the accuracy of depth estimation, the use of convex upsampling for efficient full-resolution recovery inevitably leads to the loss of high-frequency geometric details and introduces grid-like artifacts in coarse normal $\mathbf{N}_{c}$, as evident in Fig. 6. The proposed normal refinement module leverages the semantic information from radiance features and geometric cues from the coarse normal map, enabling the reconstruction of high-quality surface normals with both qualitative (Fig. 6) and quantitative (See Supp.Mat.) improvements.
@@
We introduce a feed-forward direct illumination scaling factor $s_{\text{d}}$ to compensate for the brightness loss caused by convolved environment lighting. As shown in Fig. 7, applying $s_{\text{d}}$ brings the shading closer to the ground truth. To further enhance realism by modeling shadow effects and complex light interactions such as multi-bounce illumination, we additionally learn light visibility and indirect lighting components. As illustrated in the figure, our method accurately captures occlusion-induced shadows, including those cast by the arm and ball. Across all experiments, realistic effects are achieved using only tens of light samples per Gaussian. Specifically, we sample 40 rays per Gaussian, offering a favorable trade-off between visual quality and computational efficiency. With TensorRT acceleration, our model runs at 20 FPS during inference.
@@ 5 Conclusion
This paper proposes GRGS, a generalizable and relightable 3D Gaussian framework for high-fidelity human novel view synthesis under diverse lighting conditions. GRGS adopts a supervised 2D-to-3D projection strategy to transfer geometry, material, and illumination cues from multi-view 2D observations into 3D Gaussian attributes, enabling efficient feed-forward inference without person-specific optimization. To ensure accurate and lighting-invariant geometry, we first construct a Lighting-aware Geometry Refinement module to estimate robust depth and surface normals. In addition, a Physically-Grounded Neural Rendering module is presented to integrate physics-based shading with neural inference to synthesize realistic lighting effects, avoiding the cost of explicit ray tracing. Thanks to these designs, GRGS achieves high-quality, editable human relighting and strong generalization to unseen identities and lighting environments.
@@
In this supplementary material, we provide additional information, including preliminary concepts (Sec.A), extended experimental details (Sec.B), implementation specifics (Sec.C), optimization strategies (Sec.D), and a discussion of limitations (Sec. E).
Appendix A Preliminary

### v2

@@ 3.2 Physically Grounded Neural Rendering
Based on the accurate geometry, we further propose a PGNR module to combine the generalization capabilities of neural networks with the physical accuracy of physically based rendering, enabling efficient inference of material properties and complex illumination of unseen data.
@@
To support relightable rendering, PGNR parameterizes each 3D Gaussian point $\mathcal{P}$ as a set of attributes: (1) vanilla attributes: position $\mathbf{p}\in\mathbb{R}^{3}$, rotation $\mathbf{r}\in\mathbb{R}^{4}$, scale $\mathbf{s}\in\mathbb{R}^{3}$, and opacity $\alpha\in\mathbb{R}$, (2) geometry-related attributes: surface normal $\mathbf{n}\in\mathbb{R}^{3}$ and light visibility $\mathbf{v}\in\mathbb{R}^{16}$ encoded via SH coefficients, (3) material attributes: albedo $\mathbf{a}\in\mathbb{R}^{3}$ and roughness $\gamma\in\mathbb{R}$, and (4) illumination attributes: indirect lighting $\mathbf{l}_{\text{ind}}\in\mathbb{R}^{48}$ represented by SH coefficients. We formulate these 3D attributes on corresponding 2D maps via pixel-aligned depth maps (Sec. 3.1), enabling direct inference of Gaussian point clouds without optimization. For simplicity, we consider a single-view depth map $\mathbf{D}$ and compute 3D position map $\mathbf{X}$ via unprojection (Eq. 3), while normals $\mathbf{N}_{f}$ are predicted by the refinement module. We introduce a geometry-aware encoder $\xi_{\text{geo}}$ that extracts multi-scale geometric features $\{\mathbf{F}_{s}^{g}\}_{s=1}^{K}$ from $\mathbf{D}$ and $\mathbf{N}_{f}$, encoding both 2D image features and 3D spatial geometry. These features are then fused with radiance features to incorporate appearance and geometric context for jointly modeling geometry-aware attribute maps. Specifically, we first extract the full-resolution Gaussian features $\Theta$ through a decoder $\Phi_{\text{parm}}$:
@@
where $\oplus$ denotes the concatenation operation applied across all feature levels. Subsequently, the vanilla Gaussian parameter maps (rotation map $\mathbf{R}$, scale map $\mathbf{S}$, and opacity map $\mathbf{O}$) are predicted with a Gaussian head $h_{g}$:
@@
For material attributes, we introduce a residual component $\Delta\mathbf{A}$ to accelerate the convergence of albedo estimation under diverse lighting conditions. This component acts as a delighting term, helping disentangle shadows and illumination effects from the intrinsic surface properties. We achieve this by leveraging a material prediction head $h_{m}$ to jointly estimate the residual albedo $\Delta\mathbf{A}$ and the surface roughness map $\mathbf{\Gamma}$:
@@
The final albedo is computed as: $\mathbf{A}=\text{Sigmoid}\left(\mathbf{I}+\Delta\mathbf{A}\right)$, where $\mathbf{I}$ denotes the input image. Besides, since the SH-encoded visibility map $\mathbf{Y}_{\text{vis}}$ is independent of image appearance, we predict it via another lightweight decoder $\Phi_{\text{vis}}$ conditioned solely on the geometric features $\{\mathbf{F}_{s}^{g}\}_{s=1}^{K}$.
@@
$\mathbf{L}_{\text{d}}^{\prime}\left(\boldsymbol{\omega}^{\prime}\right)=\int_{\Omega}\left(\boldsymbol{\omega}^{\prime}\cdot\boldsymbol{\omega}\right)^{l}\mathbf{L}_{\text{d}}\left(\boldsymbol{\omega}\right)d\boldsymbol{\omega}$ (10)
where $\boldsymbol{\omega}$ and $\boldsymbol{\omega}^{\prime}$ denote spherical directions in the original and convolved environment map respectively, $\Omega$ represents the unit sphere, and $l$ is the shininess exponent controlling the angular falloff in the Phong reflectance model. However, such integration may result in underexposed lighting compared to ground-truth shading, as illustrated in Sec. 4.2. To address this, we introduce a direct illumination scaling factor $s_{\text{d}}\in\mathbb{R}$ to globally compensate for the brightness discrepancy. In addition, since the SH-encoded indirect lighting map $\mathbf{Y}_{\text{ind}}$ is influenced not only by the appearance and geometry but also by the input illumination, we design a light encoder-decoder that predicts the direct illumination scaling map $\mathbf{S}_{\text{d}}$ and the indirect lighting map $\mathbf{Y}_{\text{ind}}$:
@@
To enable physically plausible light interaction on the human surface, we compute the PBR color $\mathbf{C}_{\text{PBR}}$ for each Gaussian point. Following the rendering equation [kajiya1986rendering], the outgoing radiance $\mathbf{C}_{\text{PBR}}$ in direction $\boldsymbol{\omega}_{o}$ can be given by:
$\mathbf{C}_{\text{PBR}}(\boldsymbol{\omega}_{o})=\!\!\int_{\Omega}\!\!\mathbf{L}(\boldsymbol{\omega}_{i})f(\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o},\mathbf{a},\gamma)(\boldsymbol{\omega}_{i}\!\cdot\!\boldsymbol{n})\,d\boldsymbol{\omega}_{i}$ (13)
where $f$ is the simplified Disney BRDF function [burley2012physically] modeling the surface reflectance properties, and $\Omega$ here represents the hemisphere oriented around the surface normal $\boldsymbol{n}$. $\mathbf{L}\left(\boldsymbol{\omega}_{i}\right)$ denotes the incident radiance from direction $\boldsymbol{\omega}_{i}$:
$\mathbf{L}(\boldsymbol{\omega}_{i})=\!V(\boldsymbol{\omega}_{i})\!\left(s_{\text{d}}\mathbf{L}_{\text{d}}^{\prime}(\boldsymbol{\omega}_{i})\right)+\!\left(1\!-\!V(\boldsymbol{\omega}_{i})\right)\!\mathbf{L}_{\text{ind}}(\boldsymbol{\omega}_{i})$ (14)
where $V\left(\boldsymbol{\omega}_{i}\right)$ denotes the light visibility from direction $\boldsymbol{\omega}_{i}$, parameterized using SH coefficients $\mathbf{v}$. Similarly, $\mathbf{L}_{\text{ind}}\left(\boldsymbol{\omega}_{i}\right)$ represents the indirect illumination from direction $\boldsymbol{\omega}_{i}$, also encoded via SH coefficients $\mathbf{l}_{\text{ind}}$. Although R3DGS [gao2024relightable] employs a similar lighting model, it estimates indirect illumination in an implicit manner through unsupervised inverse rendering, which often results in inaccurate predictions and limited generalization to novel lighting conditions while imposing a substantial computational burden due to the explicit ray tracing strategy. In contrast, our GRGS utilizes the strong generalization capability of a U-Net-based architecture in conjunction with a 2D-to-3D projection training strategy, enabling more accurate estimation of both visibility and indirect illumination under arbitrary lighting conditions. After computing the physically-based rendering color $\mathbf{C}_{\text{PBR}}$ for each Gaussian point, we perform rasterization to synthesize the final image.
@@
Leveraging the excellent differentiability of 3DGS [kerbl20233d], GRGS optimizes lighting parameters in the 3D space directly from 2D image supervision, thereby circumventing the costly computation of explicit ray tracing. Specifically, ground truths of ambient occlusion, direct lighting, and indirect lighting maps are used as photometric supervision signals to guide the optimization of the 3D Gaussian representation via efficient gradient-based learning. Among these signals, ambient occlusion and direct lighting maps are jointly used to provide more accurate supervision for visibility estimation. While ambient occlusion captures geometric occlusion effects, direct lighting offers complementary shadow-related illumination cues, and their combination helps enforce visibility consistency between 2D renderings and the underlying 3D geometry. For effective indirect lighting learning within the rendering equation, we employ a gradient-truncated hard shadow fusion scheme, which truncates gradients through visibility to prevent erroneous gradient propagation and leverages hard shadow fusion to accelerate the convergence of indirect lighting.
@@ 4 Experiments
We utilize two human scan datasets, Twindom [twindom] and THuman2.0 [yu2021function4d], to validate the effectiveness of our method. To ensure high-quality rendering, we carefully selected 800 scans from Twindom and 337 scans from THuman2.0, filtering out meshes with noticeable artifacts. Additionally, 384 HDR environment maps were collected from Polyhaven, HDRMAPS, and iHDRI to simulate diverse illumination conditions. More dataset construction details are listed in Supp.Mat.. In addition, we also employ multi-view real-world data from GPS-Gaussian [zheng2024gps] and DNA-Rendering [cheng2023dna] to further evaluate the generalization ability of our method in real scenarios.
Table 1: Quantitative comparison on the synthetic dataset. In accordance with [zheng2024gps], the SSIM and LPIPS metrics are calculated within the bounding box delineating the human region, while the MAE of normal and PSNR metrics is computed within the foreground mask.
@@
Figure 5: Qualitative comparison between our method and 2D image-based approaches. The bottom row shows results from the real dataset [zheng2024gps]. Zoom in for the best view.
@@
We compare our method on our synthetic dataset and real dataset [zheng2024gps] against two 3DGS-based approaches, R3DGS [gao2024relightable] and ARGS [li2024animatable], as well as two 2D image-based methods, IC-Light [zhang2025scaling] and SwitchLight [kim2024switchlight]. Note that since 2D methods cannot synthesize novel views, we directly use ground-truth novel view images as their input for a fair comparison. For quantitative evaluation, we employ PSNR, SSIM, and LPIPS to evaluate the diffuse albedo, ambient occlusion, and relighting quality. Further, we utilize MAE to evaluate the normal map.
@@ 4.1 Comparison Results
We first make a comparison with 3D-based methods (i.e., R3DGS [gao2024relightable] and ARGS [li2024animatable]) from the perspective of material estimation, geometry reconstruction, and relighting quality, as shown in Fig. 4 and Table 1. R3DGS [gao2024relightable] struggles to recover smooth, detailed normals from sparse views without human-specific priors, leading to suboptimal relighting performance. ARGS [li2024animatable] alleviates this by introducing a body template prior. However, optimizing geometry solely from rendered images leads to overly smooth surfaces lacking high-frequency details. Moreover, both methods depend on inverse rendering under an inherently ill-posed formulation, which yields inaccurate material estimation and consequently degrades the overall quality of relighting results. Thanks to the proposed LGR and PGNR modules, GRGS produces high-quality geometry and realistic relighting results under novel viewpoints.
Furthermore, a comparison with representative 2D-based relighting methods (i.e., IC-Light [zhang2025scaling] and SwitchLight [kim2024switchlight]) is illustrated in Fig. 5. The baseline methods produce noticeable shading and lighting variations that often lack physical plausibility, exhibiting artifacts such as inconsistent light–geometry interactions and unnatural shadow orientations, thereby diminishing overall realism. In contrast, our results closely match the ground-truth images generated by a physically-based renderer and deliver superior perceptual quality. Additional evidence is provided by a user study (see Supp.Mat.).
@@ 4.2 Ablation Study
Geometry reconstruction.
A major limitation of the original GPS-Gaussian [zheng2024gps] is that it struggles to maintain robustness against uneven lighting on the human surface without lighting priors, leading to a significant degradation in geometric consistency, as shown in Fig. 3. In contrast, our LGR module learns lighting-robust radiance features under diverse illumination conditions, thereby improving reconstruction accuracy under uneven lighting and enhancing both geometric consistency and NVS quality (Table. 2).
Although the lighting-robust radiance features enhance depth estimation accuracy, the use of convex upsampling for full-resolution recovery inevitably causes the loss of high-frequency geometric details and introduces checkerboard artifacts in the coarse normal map $\mathbf{N}_{c}$, as shown in Fig. 6. The proposed normal refinement module leverages both semantic information from radiance features and geometric cues from the coarse normal map to reconstruct high-quality surface normals with consistent improvements in qualitative (Fig. 6) and quantitative (see Supp.Mat.) results.
Table 2: NVS and Depth Evaluation.
Method NVS Depth
[…]
GPS-Gaussian [zheng2024gps] 30.444 0.913 0.107 1.463 68.91
Ours 31.694 0.949 0.067 0.692 85.39
@@
To mitigate the brightness loss caused by convolved environment lighting, we introduce a feed-forward direct illumination scaling factor $s_{\text{d}}$, which yields shading closer to the ground truth as shown in Fig. 7. To further enhance realism by modeling shadow effects and complex light interactions such as multi-bounce illumination, we additionally learn light visibility and indirect lighting. As illustrated in the figure, our method accurately captures occlusion shadows from the arm and ball. Across all experiments, realistic effects are achieved with only tens of light samples per Gaussian. Specifically, sampling 40 rays strikes an effective balance between quality and efficiency. With TensorRT acceleration, our model runs at 20 FPS during inference. Further analysis of computational efficiency is provided in Supp.Mat..
@@ 5 Conclusion
This paper proposes GRGS, a generalizable and relightable 3D Gaussian framework for high-fidelity human novel view synthesis under diverse lighting conditions. GRGS adopts a supervised 2D-to-3D projection strategy to transfer geometry, material, and illumination cues from multi-view 2D observations into 3D Gaussian attributes, enabling efficient feed-forward inference without person-specific optimization. To ensure accurate geometry, we first construct a Lighting-robust Geometry Refinement module to estimate robust depth and surface normals. In addition, a Physically Grounded Neural Rendering module is presented to integrate physics-based shading with neural inference to synthesize realistic lighting effects, avoiding the cost of explicit ray tracing. Thanks to these designs, GRGS achieves high-quality, editable human relighting and strong generalization to unseen identities and lighting environments.
[…]
@@
In this supplementary material, we provide additional information, including preliminary concepts (Sec.6), implementation details (Sec.7), more comparison results (Sec.8), more analysis on core component (Sec.9), detailed loss functions(Sec.10) and a discussion of limitations (Sec. 11).
6 Preliminary